QUESTION 1 Florida Mining Equipment Depot conducts recapitalization (or overhaul) operations on large, heavy mining equipment. Vehicles scheduled for periodic overhaul and repair are brought to the depot and a complete overhaul returns them back to the minim company in a like new condition. In order for ope smoothly, spare parts must be available so that the production line does not idle. Below is a table of three critical parts used in the overhaul process. The annual holding and backorder costs are given below. The backorder cost for each item is a one-time backorder cost (the cost does not accrue over time) and represents the cost of expediting that specific part (backorder costs may differ from item to item based upon the specific characteristics of the item and the expediting procedures). The holding cost is the cost to run ot holding one unit of that particular part number in inventory for one year. Assume 52 weeks per year If the optimal order quantity (economic order quantity) for part 1002 is 60 what will be the holding cost savings if the order quantity is cut in halt? Avg Std Dev Current On- Annual Unit Weekly Weekly Lead Time Hand Ordering Holding Backorder Part Cost Dmd Dmd (in weeks) Inventory Cost Cost Cost 1001 $8.50 19 6 1002 $47.25 75 38 $125 $55 $1,000 112 $225 $45 $1,000 3 $175 $125 $1,000 16 1003 $202.50 0.5 0.8 $682 $437 $785 O $251
Answer #1

For 1003

Average Weekly Demand, d = 0.5 unit

Annual Demand = 0.5*52 = 26 unit

Current inventory on hand = 3

So Actual Annual Demand, D = 26-3 = 23 unit

Order cost, S = 175

Annual Holding Cost, IC = 125

Optimal Order quantity , EOQ, Q = sqrt(2*D*S/IC) =sqrt(2*23*175/125) = sqrt(64.4) = 8.02 = 8

Answer = 8